# Text-to-Image Rectified Flow as Plug-and-Play Priors

### v2

@@ Abstract
Large-scale diffusion models have achieved remarkable performance in generative tasks. Beyond their initial training applications, these models have proven their ability to function as versatile plug-and-play priors. For instance, 2D diffusion models can serve as loss functions to optimize 3D implicit models. Rectified flow, a novel class of generative models, enforces a linear progression from the source to the target distribution and has demonstrated superior performance across various domains. Compared to diffusion-based methods, rectified flow approaches surpass in terms of generation quality and efficiency, requiring fewer inference steps. In this work, we present theoretical and experimental evidence demonstrating that rectified flow based methods offer similar functionalities to diffusion models — they can also serve as effective priors. Besides the generative capabilities of diffusion priors, motivated by the unique time-symmetry properties of rectified flow models, a variant of our method can additionally perform image inversion. Experimentally, our rectified flow-based priors outperform their diffusion counterparts — the SDS and VSD losses — in text-to-3D generation. Our method also displays competitive performance in image inversion and editing. Code is available at: https://github.com/yangxiaofeng/rectified_flow_prior.
@@ 1 Introduction
Large-scale diffusion models [1; 2; 3; 4] have made significant advances in recent years. Applications in image generation [4; 5; 6; 7; 8], video creation [9; 10], and music production [11] have shown superior performance compared to traditional Generative Adversarial Networks (GANs) [12]. In addition to excelling in tasks for which they are specifically trained, various studies [13; 14; 15; 16] indicate that these models can also serve as plug-and-play priors for related tasks, i.e., using the pretrained diffusion models as loss functions. For example, Dreamfusion [14] introduces the SDS loss, utilizing a pretrained text-to-image diffusion model to generate 3D objects. DDS [17] extends the original SDS loss to address image editing challenges. Various subsequent works [15; 16; 18; 19] further refine the SDS loss to enhance generation quality and diversity.
@@
Rectified flow [20; 21; 22] is a new type of flow-based generative model that linearly transfers the source distribution to the target distribution. This property of linear transfer facilitates simulation-free training and faster sampling at inference time. Recently, methods based on rectified flow have garnered considerable research interest. Initial work such as InstalFlow [23] has been capable of generating high-quality images in just a few inference steps. More recent developments, like rectified flow based stable diffusion [24], have achieved new state-of-the-art performance in image generation. Given the growing research focus, it remains to be seen whether large-scale pretrained rectified flow models can function as effective priors, similar to diffusion-based methods.
@@
In this work, we explore how to distill knowledge from pretrained text-to-image rectified flows and use them as priors for other tasks. We propose three methods. The first, named RFDS (Rectified Flow Distillation Sampling), is intended to serve as a baseline method for rectified flow priors, analogous to the SDS loss in diffusion models. Specifically, we derive the baseline RFDS by calculating the gradients of the input image by reversing the rectified flow training process. We observe a similar phenomenon to that seen in diffusion-based methods: by removing the Jacobian of the rectified flow U-Net, RFDS can generate meaningful images or 3D objects given text conditions.
@@
Moreover, an intriguing characteristic of rectified flow is its linear trajectory between the initial Gaussian distribution and the image distribution, a property known as the time-symmetry property. Inspired by this characteristic, we notice that the baseline RFDS can be easily extended to optimize randomly sampled noise by taking the negative gradient of the original RFDS. We refer to this method as iRFDS, which represents the inverse of RFDS. Optimizing noise is crucial for tasks like image inversion and editing. Typically, an image editing problem involves identifying the original noise of a given real image and subsequently altering the text prompt to edit the image. Our experiments demonstrate that iRFDS is effective at performing image inversion and editing on real-world images.
@@
Finally, we introduce RFDS-Rev (RFDS Reversal) to enhance the generation quality of the baseline RFDS loss. Our observations reveal that the original RFDS loss exhibits behaviors similar to those of the SDS loss, such as limited detail and diversity in generated content. We attribute these issues to the mode-seeking tendency of the RFDS loss due to random sampling. As rectified flow learns a unique mapping for each noise input, the RFDS leads to averaged images with diminished detail when random noise sampling is used. To address these challenges, we propose RFDS-Rev, which iteratively applies the iRFDS for flow reversal to determine the original noise, and RFDS for knowledge distillation to refine the input. Experimentally, we find that the RFDS-Rev method significantly outperforms both the RFDS baseline and the diffusion-based VSD loss.
@@
Our experiments are carried out on 2D image inversion and editing for iRFDS and text-to-3D generation for RFDS and RFDS-Rev. Notably, the RFDS-Rev method achieves a new state-of-the-art performance in text-to-3D benchmarks among all 2D lifting methods, surpassing a broad spectrum of diffusion priors. Some generation results can be found in Fig. 1.
@@
2 Preliminaries: Rectified Flow
The idea of rectified flow is proposed in multiple concurrent works [20; 21; 22]. Here, we follow the interpretation of  [20] to briefly introduce the basic concepts of rectified flow.
Rectified flow is an ordinary differentiable equation model that defines the transportation from $\pi_{0}$ to $\pi_{1}$ to follow a straight line. Its general form is defined as:
$\text{d}Z_{t}=v(Z_{t},t)\text{d}t.\;$ (1)
Given two data points sampled from the two distributions: $X_{0}\sim\pi_{0}$, $X_{1}\sim\pi_{1}$. The drift force $v$ is set to learn the difference between the two data points:
$\underset{\phi}{\text{min}}\;\int_{0}^{1}\mathbb{E}\left[\parallel(X_{1}-X_{0})-v_{\phi}(X_{t},t)\parallel^{2}\right]\text{d}t,\;\text{with}\;X_{t}=tX_{1}+(1-t)X_{0}.$ (2)
Practically, the drift force $v$ is parameterized using neural networks with parameters $\phi$. Once the network is trained with the above loss function, data points in the target distribution can be recovered by following the ODE process in just a few steps along straight flow lines.
However, it is important to note that the primary objective of this work is not to recover the original data points (e.g., generating images using rectified flow networks). Instead, our aim is to distill knowledge from such pretrained networks —specifically by employing the pretrained network as a loss function—to enhance the functionality of text-to-image networks across a wider array of applications, including image editing and text-to-3D generation.
Our proposed method is also related to diffusion-based priors. We discuss these related works in the Appendix.
@@ 3 Methods
Formulation. The problem of optimization using pretrained rectified flow models as loss functions can be formatted as follows: we would like to optimize $x=g(\theta)$ given $\theta^{\ast}=\text{argmin}_{\theta}\mathcal{L}_{\text{Rf}}(\phi,x=g(\theta))$. The form of $x$ can be Neural Radiance Field or 3D Gaussian Splitting in 3D generation problem, where the parameter $\theta$ denotes the parameters of the 3D models. In the 2D case, $x$ is directly $\theta$. In the remaining of this section, we prove how to define the $\mathcal{L}_{\text{Rf}}$ and find its gradients related to $\theta$.
@@ 3.1 RFDS
Consider the training loss function of rectified flow in Eq. 2. Using the pretrained rectified flow network as a loss function can be seen as an inverse operation of the training process; that is, optimizing the input rather than the model parameters. Specifically, when we place the optimization variable $x$ in the aforementioned equation, it transforms as follows:
[…]
Here, the rectified flow network parameter $\phi$ is fixed and $\epsilon$ is a randomly sampled noise.
To find $\theta$, the gradients of the above equation with respect to $\theta$ can be written as:
$\nabla_{\theta}\mathcal{L}_{\text{Rf}}(\phi,x,\epsilon,t)=2\times\mathbb{E}\left[\underbrace{(x-\epsilon-v_{\phi}(x_{t},t))}_{\text{Rf Residual}}\frac{\partial(x-\epsilon-v_{\phi}(x_{t},t))}{\partial x}\underbrace{\frac{\partial x}{\partial\theta}}_{\text{Generator Jacobian}}\right].$ (4)
It could be further simplified as:
$\nabla_{\theta}\mathcal{L}_{\text{Rf}}(\phi,x,\epsilon,t)=2\times\mathbb{E}\left[\underbrace{(x-\epsilon-v_{\phi}(x_{t},t))}_{\text{Rf Residual}}(\underbrace{\frac{\partial(x-\epsilon)}{\partial x}}_{\text{Positive Constant}}-\underbrace{\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}}_{\text{Rf network Jacobian}}\frac{\partial x_{t}}{\partial x})\underbrace{\frac{\partial x}{\partial\theta}}_{\text{Generator Jacobian}}\right].$ (5)
As highlighted in diffusion-based methods [14], calculating the network Jacobian is computationally expensive and reacts inadequately to low levels of noise. We observe similar behaviors in models based on rectified flow. Consequently, we also choose to disregard the rectified flow network Jacobian by setting it to identity matrix, akin to the approach taken with diffusion priors. This modification leads to the final RFDS loss:
$\nabla_{\theta}\mathcal{L}_{\text{Rf}}(\phi,x,\epsilon,t)\simeq\mathbb{E}\left[w(t)\underbrace{(x-\epsilon-v_{\phi}(x_{t},t))}_{\text{Rf Residual}}\underbrace{\frac{\partial x}{\partial\theta}}_{\text{Generator Jacobian}}\right].$ (6)
Here we use $w(t)$ to represent a positive value related to timestep $t\in[0,1]$, and it absorbs all constant values in Eq. 5. The RFDS loss is able to use the pretrained rectified flow model as a loss function to optimize any given image $x$. Detailed algorithm of RFDS is included in the Appendix.
@@
A particularly useful and intriguing characteristic of rectified flow models is their time-symmetry property [20]. This allows the models to flow from noise to image and also from image back to noise. Motivated by this property, we observe that if our optimization objective changes from the input image $x=g(\theta)$ to the noise $\epsilon$, Eq. 5 now becomes:
$\nabla_{\epsilon}\mathcal{L}_{\text{Rf}}(\phi,x,\epsilon,t)=2\times\mathbb{E}\left[\underbrace{(x-\epsilon-v_{\phi}(x_{t},t))}_{\text{Rf Residual}}(\underbrace{\frac{\partial(x-\epsilon)}{\partial\epsilon}}_{\text{Negative Constant}}-\underbrace{\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}}_{\text{Rf network Jacobian}}\frac{\partial x_{t}}{\partial\epsilon})\right].$ (7)
Since the gradient is now calculated with respect to $\epsilon$, the term of $\frac{\partial(x-\epsilon)}{\partial\epsilon}$ will become negative. Therefore, the final form of iRFDS can be represented as:
$\nabla_{\epsilon}\mathcal{L}_{\text{Rf}}(\phi,x,\epsilon,t)\simeq\mathbb{E}\left[-w(t)\underbrace{(x-\epsilon-v_{\phi}(x_{t},t))}_{\text{Rf Residual}}\right].$ (8)
Optimizing and recovering the original noise from an image is crucial for image inversion and editing problems. By first retrieving the original noise and subsequently modifying the caption, we demonstrate in the experiment section that our method can effectively edit real-world images. Detailed algorithm of iRFDS is included in the Appendix.

### v3

@@ Abstract
Large-scale diffusion models have achieved remarkable performance in generative tasks. Beyond their initial training applications, these models have proven their ability to function as versatile plug-and-play priors. For instance, 2D diffusion models can serve as loss functions to optimize 3D implicit models. Rectified Flow, a novel class of generative models, has demonstrated superior performance across various domains. Compared to diffusion-based methods, rectified flow approaches surpass them in terms of generation quality and efficiency. In this work, we present theoretical and experimental evidence demonstrating that rectified flow based methods offer similar functionalities to diffusion models — they can also serve as effective priors. Besides the generative capabilities of diffusion priors, motivated by the unique time-symmetry properties of rectified flow models, a variant of our method can additionally perform image inversion. Experimentally, our rectified flow based priors outperform their diffusion counterparts — the SDS and VSD losses — in text-to-3D generation. Our method also displays competitive performance in image inversion and editing. Code is available at: https://github.com/yangxiaofeng/rectified_flow_prior.
@@ 1 Introduction
Large-scale diffusion models (Ho et al., 2020; Song & Ermon, 2019; 2020; Rombach et al., 2022; Karras et al., 2022) have made significant advances in recent years. Applications in image generation (Rombach et al., 2022; Dhariwal & Nichol, 2021; Ho et al., 2022; Ramesh et al., 2022; Saharia et al., 2022), video creation (Guo et al., 2023b; Blattmann et al., 2023), and music production (Huang et al., 2023) have demonstrated superior performance compared to traditional Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Brock, 2018). Beyond excelling in tasks for which they are specifically trained, various studies (Graikos et al., 2022; Poole et al., 2022; Wang et al., 2023; Yu et al., 2024) indicate that these models can also serve as plug-and-play priors for related tasks, i.e., using the pretrained diffusion models as loss functions. For instance, Dreamfusion (Poole et al., 2022) introduces the SDS loss, utilizing a pretrained text-to-image diffusion model to generate 3D objects. DDS (Hertz et al., 2023) extends the original SDS loss to address image editing challenges. Several subsequent works (Wang et al., 2023; Yu et al., 2024; Katzir et al., 2024; Yang et al., 2023) have refined the SDS loss to further enhance generation quality and diversity, applying diffusion priors to a variety of downstream applications, including 3D editing, 4D generation, and human avatar creation.
@@
Rectified flow (Liu et al., 2022; Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022) is a special type of flow-based method that connects data and noise on a straight line. Recently, methods based on rectified flow have garnered considerable research interest. Initial work such as InstalFlow (Liu et al., 2024) has demonstrated the capability to generate high-quality images in just a few inference steps. More recent developments, including SiT (Ma et al., 2024) and Stable Diffusion 3 (Esser et al., 2024), have achieved new state-of-the-art performance in image generation. Given this growing research focus, it remains to be seen whether large-scale pretrained rectified flow models can function as effective priors, similar to diffusion-based methods.
@@
In this work, we explore how to distill knowledge from pretrained text-to-image rectified flow models and use them as priors for other tasks. To ensure our approach is generalizable and applicable to a wide range of flow-matching methods, we develop our method based on the generalized Stochastic Interpolants framework (Albergo et al., 2023) rather than directly deriving it from the rectified flow formulation. We then incorporate specific interpolation factors of rectified flow to adapt our method to this particular model type. We propose three methods. The first, named RFDS (Rectified Flow Distillation Sampling), is intended to serve as a baseline method for rectified flow priors, analogous to the SDS loss in diffusion models. Specifically, we derive the baseline RFDS by calculating the gradients of the input image by reversing the flow-matching training process. We observe a phenomenon similar to that seen in diffusion-based methods: by removing the Jacobian of the rectified flow network, RFDS can generate meaningful images or 3D objects given text conditions.
@@
Moreover, inspired by the straight line trajectories of rectified flow, we notice that the baseline RFDS can be easily extended to optimize randomly sampled input noise by taking the “negative gradient” of the original RFDS. We refer to this method as iRFDS, which represents the inverse of RFDS. Optimizing noise is crucial for tasks like image inversion and editing. Typically, an image editing problem involves identifying the original noise of a given real image and subsequently altering the text prompt to edit the image. Our experiments demonstrate that iRFDS is effective at performing image inversion and editing on real-world images.
@@
Finally, we introduce RFDS-Rev (RFDS Reversal) to enhance the generation quality of the baseline RFDS loss. Our observations reveal that the original RFDS loss exhibits behaviors similar to those of the SDS loss, such as limited detail and diversity in generated content. We attribute these issues to the mode-seeking tendency of the RFDS loss due to random sampling. The baseline RFDS leads to averaged images with diminished detail when random noise sampling is used. To address these challenges, we propose RFDS-Rev, which iteratively applies iRFDS for flow reversal to determine the original noise, and RFDS for knowledge distillation to refine the input. Experimentally, we find that the RFDS-Rev method significantly outperforms both the RFDS baseline and the diffusion-based VSD loss.
@@
Additionally, we demonstrate that our method is applicable to score-matching methods, such as diffusion models. The diffusion models, trained with score-matching objectives, can be written in deterministic PF-ODEs, making them compatible with our proposed methods. Notably, when we formulate score-matching models in terms of PF-ODEs, our proposed RFDS baseline becomes identical to the SDS loss. Furthermore, our introduced iRFDS and RFDS-Rev expand upon previous SDS-like, score-matching-based knowledge distillation approaches, broadening their scope and applicability.
Our experiments are conducted on 2D image inversion and editing for iRFDS and text-to-3D generation for RFDS and RFDS-Rev. Notably, the RFDS-Rev method achieves a new state-of-the-art performance in text-to-3D benchmarks among all 2D lifting methods, surpassing a broad spectrum of diffusion priors. Some generation results can be found in Fig. 1.
@@
2 Preliminaries: Flow-based models and Rectified Flow
The idea of rectified flow is proposed in multiple concurrent works (Liu et al., 2022; Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022). Here, we follow the interpretation of Stochastic Interpolants (Albergo & Vanden-Eijnden, 2022; Ma et al., 2024) to briefly introduce the basic concepts.
The Stochastic Interpolants framework starts by defining a general process of interpolating noise and image:
$\bm{x_{t}}=\alpha_{t}\bm{\bm{x_{*}}}+\sigma_{t}\bm{\bm{\epsilon}}.\;$ (1)
Here, $\bm{x_{*}}\sim p(\bm{x})$ represents the data sampled from the data distribution, while $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ denotes noise drawn from a standard Gaussian distribution. $\alpha_{t}$ and $\sigma_{t}$ are two pre-defined variables used to control the interpolation trajectory and are only related to $t$. Rectified flow is a special case of the interpolation in Eq. 1 by defining $\alpha_{t}$ and $\sigma_{t}$ to change linearly with time $t$. In conditional flow matching (Lipman et al., 2022), the parameters are set as $\alpha_{t}=1-t$ and $\sigma_{t}=t$. In rectified flow (Liu et al., 2022), the settings are reversed, with $\alpha_{t}=t$ and $\sigma_{t}=1-t$.
The process of obtaining $\bm{x_{t}}$ can also be formulated as a probability flow ODE (PF-ODE) with a velocity field:
$\dot{\bm{x_{t}}}=\bm{v}(\bm{x_{t}},t),\;$ (2)
$\bm{v}(\bm{x_{t}},t)=\dot{\alpha_{t}}\mathop{\mathbb{E}}[\bm{x_{*}}\mid\bm{x_{t}}=\bm{x}]+\dot{\sigma_{t}}\mathop{\mathbb{E}}[\bm{\epsilon}\mid\bm{x_{t}}=\bm{x}],\;$ (3)
where $\dot{\alpha_{t}}$ and $\dot{\sigma_{t}}$ are the gradients of $\alpha_{t}$ and $\sigma_{t}$ with respect to $t$. We refer readers to Stochastic Interpolants (Albergo et al., 2023) and SiT (Ma et al., 2024) for detailed proofs of the above two equations.
Practically, flow-based generative models learn the velocity $\bm{v}$ by parameterizing it with neural networks using parameters $\bm{\phi}$ and optimizing a flow-matching objective:
$\underset{\bm{\phi}}{\text{min}}\;\int_{0}^{1}\mathbb{E}\left[\parallel\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x_{*}}-\dot{\sigma_{t}}\bm{\epsilon}\parallel^{2}\right]\text{d}t.\;$ (4)
However, it is important to note that the primary objective of this work is not to sample $\bm{x_{t}}$ (e.g., generating images using rectified flow networks). Instead, our aim is to distill knowledge from such pretrained networks—specifically by employing the pretrained network as a loss function—to enhance the functionality of image generation networks across a wider array of applications, including image editing and text-to-3D generation.
Our proposed method is also related to widely studied diffusion-based priors. We discuss these related works in Appendix Sec C.
@@ 3 Methods
Formulation. The problem of optimization using pretrained rectified flow models $\bm{v_{\phi}}$ as loss functions can be formatted as follows: we would like to optimize $\bm{x}=g(\bm{\theta})$ by $\bm{\theta^{\ast}}=\text{argmin}_{\bm{\theta}}\mathcal{L}(\bm{\phi},\bm{x}=g(\bm{\theta}))$. The form of $\bm{x}$ can be Neural Radiance Field or 3D Gaussian Splatting (Kerbl et al., 2023) in the 3D generation problem, where the parameter $\bm{\theta}$ denotes the parameters of the 3D models. In the 2D case, $\bm{x}$ is directly $\bm{\theta}$. In the remainder of this section, we prove how to define the loss $\mathcal{L}$ and find its gradients with respect to $\bm{\theta}$.
@@ 3.1 RFDS
Consider the training loss function of flow-based generative models in Eq. 4. Using the pretrained network as a loss function can be seen as optimizing the input rather than the model parameters. Specifically, when we place the optimization variable $\bm{x}$ in the aforementioned equation, it transforms as follows:
$\int_{0}^{1}\mathbb{E}\left[\parallel\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon}\parallel^{2}\right]\text{d}t,\;\text{with}\;\bm{x}=g(\bm{\theta})\;\text{and}\;\bm{x_{t}}=\alpha_{t}\bm{x}+\sigma_{t}\bm{\epsilon}.$ (5)
Here, the network parameter $\bm{\phi}$ is fixed and $\bm{\epsilon}$ is a randomly sampled noise.
To find $\bm{\theta}$, the gradients of the above equation with respect to $\bm{\theta}$ can be written as:
$\nabla_{\bm{\theta}}\mathcal{L}_{\text{rfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)=2\times\mathbb{E}\left[\underbrace{(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}_{\text{Flow Residual}}(\frac{\partial(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}{\partial\bm{x}})\underbrace{\frac{\partial\bm{x}}{\partial\bm{\theta}}}_{\text{Generator Jacobian}}\right].$ (6)
It can be further simplified as:
$\nabla_{\bm{\theta}}\mathcal{L}_{\text{rfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)=2\times\mathbb{E}\left[\underbrace{(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}_{\text{Flow Residual}}(\underbrace{\frac{\partial(-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}{\partial\bm{x}}}_{-\dot{\alpha_{t}}}+\underbrace{\frac{\partial\bm{v_{\phi}}(\bm{x_{t}},t)}{\partial\bm{x_{t}}}}_{\text{Network Jacobian}}\underbrace{\frac{\partial\bm{x_{t}}}{\partial\bm{x}}}_{\alpha_{t}})\underbrace{\frac{\partial\bm{x}}{\partial\bm{\theta}}}_{\text{Generator Jacobian}}\right].$ (7)
As highlighted in diffusion-based methods (Poole et al., 2022), calculating the network Jacobian is computationally expensive and reacts inadequately to low levels of noise. We observe similar behaviors in models based on flow-matching. Consequently, we also choose to disregard the rectified flow network Jacobian by setting it to the identity matrix, akin to the approach taken with diffusion priors. This modification leads to the final RFDS loss:
$\nabla_{\bm{\theta}}\mathcal{L}_{\text{rfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)\simeq\mathbb{E}\left[w(t)\underbrace{(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}_{\text{Flow Residual}}\underbrace{\frac{\partial\bm{x}}{\partial\bm{\theta}}}_{\text{Generator Jacobian}}\right].$ (8)
Here we use $w(t)$ to represent a value related to timestep $t$, and it absorbs all constant values in Eq. 7. The RFDS loss is able to use the pretrained rectified flow model as a loss function to optimize a given image $\bm{x}$. Given a rectified flow based model, we can easily obtain its corresponding RFDS loss by setting $\dot{\alpha_{t}}$ and $\dot{\sigma_{t}}$ to $-1$ and $1$ or vice versa, depending on the way it formulates Eq. 1. We include a detailed algorithm for optimization using RFDS is in Appendix Sec D.
@@
A particularly interesting characteristic of the velocity prediction objective is its time-symmetry property (Liu et al., 2022). This allows the models to flow from noise to image and also from image back to noise. Motivated by this property, we observe that our method can be easily extended to optimize the noise. If our optimization objective changes from the input image $\bm{x}=g(\bm{\theta})$ to the noise $\bm{\epsilon}$, Eq. 7 now becomes:
$\nabla_{\bm{\epsilon}}\mathcal{L}_{\text{irfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)=2\times\mathbb{E}\left[\underbrace{(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}_{\text{Flow Residual}}(\underbrace{\frac{\partial(-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}{\partial\bm{\epsilon}}}_{-\dot{\sigma_{t}}}+\underbrace{\frac{\partial\bm{v_{\phi}}(\bm{x_{t}},t)}{\partial\bm{x_{t}}}}_{\text{Network Jacobian}}\underbrace{\frac{\partial\bm{x_{t}}}{\partial\bm{\epsilon}}}_{\sigma_{t}})\right].$ (9)
Therefore, the final form of iRFDS can be represented as:
$\nabla_{\bm{\epsilon}}\mathcal{L}_{\text{irfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)\simeq\mathbb{E}\left[w^{\prime}(t)\underbrace{(\bm{v_{\phi}}(\bm{x_{t}},t)-\dot{\alpha_{t}}\bm{x}-\dot{\sigma_{t}}\bm{\epsilon})}_{\text{Flow Residual}}\right],$ (10)
where the value $w^{\prime}(t)$ should have opposite signs to $w(t)$ due to the relations between $\sigma_{t}$ and $\alpha_{t}$ as stated in Sec. 2.
Optimizing and recovering the original noise from an image is crucial for image inversion and editing problems. By first retrieving the original noise and subsequently modifying the caption, we demonstrate in the experiment section that our method can effectively edit real-world images. A detailed algorithm of iRFDS is included in Appendix Sec D.